OpenStudy (anonymous):

what is the length of a prism whit a volume of 600 in 3, a length that is twice its width, and a height of 12in? (V=1x w x h)

OpenStudy (anonymous):

volume = length * width * height length = 2 width in width = width in height = 12 in 600 in^3 = (2w in)(w in)(12 in) solve for w. 50 = 2w^2 sqrt(25) = w 5 = w length = 2w = 2*5 = 10 in width = 5 in height = 12 in. Check: v = lwh... 600=10*5*12? ... CHECK!